# Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning

### v2

@@ Appendix B Mechanistic Analysis and Policy Interpretability
We also analyze attention patterns conditioned on linguistic structures. Constraints involving coordination receive 1.8 higher attention than those involving simple modification, and the policy learns to delay propagation of constraints in garden-path regions until sufficient context is available, mimicking human parsing strategies (lewis2014ccg).
On sentences with nested relative clauses (e.g., “The manager [who hired the employee [who left]] resigned”), the policy discovers a middle-out propagation strategy. Rather than processing constraints bottom-up (inner clauses first) or top-down (outer attachment first), the policy establishes clause boundaries at mid-depth first, reducing backtracks from 2.3 to 1.5 on average compared to standard heuristics. This strategy emerges without explicit supervision, suggesting the policy discovers structural regularities through reward optimization alone shihab2025detecting. The middle-out strategy processes constraints at depths 2–3 first (relative clause boundaries), establishing structural anchors before resolving attachment ambiguities at depths 1 and 4+, which reduces backtracking by 34% compared to bottom-up processing. A visualization of this depth-ordered propagation pattern is provided in the supplementary materials.
@@ B.1 Linguistic Error Analysis
We characterize error cases by sentence length, ambiguity, constraint density, and schema depth. Errors concentrate on sentences with high ambiguity (coordination, attachment) and long-distance dependencies. For transfer stress tests, we evaluate hard transfer directions: TSP to Scheduling achieves only 74% of specialist performance with 5–10 steps, but recovers to 87% with 20–30 gradient steps or a small adapter layer (2-layer MLP, 5K parameters). To test robustness to determinism violations, we simulate stochastic propagators (randomized pruning order with 10% noise) and find that the policy maintains 92% of deterministic performance, with entropy-aware fallback shihab2025differentiable providing additional robustness.
@@ Appendix C Implementation Details
This appendix provides additional implementation details for reproducibility. The MetaJuLS solver is implemented in Python 3.8 using PyTorch 1.12 for the neural network components (lecun2015deep; goodfellow2016deep) and the OR-Tools library for the base CP solver functionality (perron2008or; schulte2010gecode). The GAT architecture uses the PyTorch Geometric library for efficient graph operations (velickovic2018graph; kipf2017semi).
The training procedure uses a distributed setup with 8 GPU workers, each generating episodes in parallel (schulman2017proximal; lillicrap2015continuous). For each training round, workers collect a fixed number of episodes, which are then aggregated and used for several epochs of PPO updates before being discarded, following the standard on-policy protocol without long-term replay (mnih2015human; silver2016alphago).
@@ C.1 Hyperparameters and Robustness
| Optimizer | Adam (kingma2014adam) |
@@
Table 8 details the full model configuration. The reward function hyperparameters are set to and based on preliminary experiments that balanced domain reduction and the deterministic cost proxy described in the main text (sutton2018reinforcement; schulman2017proximal). The discount factor encourages long-term planning while preventing excessive focus on distant future rewards (schulman2017proximal; silver2017alphago).
@@ Appendix D Constraint Programming Experiments: Extended Results
Having established MetaJuLS’s effectiveness on linguistic constraints, we test universality: do the learned scheduling principles transfer to non-linguistic constraint satisfaction? This bidirectional transfer, where policies learned on CP benchmarks also improve NLP inference and vice versa, would validate that MetaJuLS captures domain-general propagation structure rather than linguistic priors. We evaluate on classical constraint programming benchmarks from MiniZinc and XCSP (van2011minizinc; lecoutre2011xcsp), training on 500 Knapsack, 300 TSP, and 200 Graph Coloring instances. We compare against JuLS (FIFO), random scheduling, dom/wdeg heuristics, Google OR-Tools (perron2008or), and Activity-Based adaptive heuristics (marques2012activity), using identical time limits (1200s per instance) and hardware.
@@
On the XCSP Competition dataset, which focuses on high-difficulty instances (lecoutre2011xcsp), the performance advantages of MetaJuLS are even more pronounced. The learned policies excel at identifying bottleneck constraints in complex constraint graphs (velickovic2018graph), leading to more effective search space pruning (bessiere2006constraint). We observe similar trends in optimality gaps and runtime improvements, with MetaJuLS achieving approximately 1.5 speedup over OR-Tools on average (perron2008or).
@@ D.1 Policy Analysis and Transfer
The learned policies reveal both rediscovery of classical heuristics and novel strategies. On Knapsack instances, the policy correlates strongly () with the Domain-over-Weight heuristic (apt2003principles), validating that learning recovers known effective strategies. On Graph Coloring, the policy learns to focus on vertices with high violation potentials rather than degree alone, outperforming classical heuristics on dense graphs (regin1994filtering; gent2006watched). Policies trained on one domain transfer effectively: Knapsack-trained policies achieve 82% of specialist performance on Bin Packing, TSP-trained achieve 74% on Scheduling, and mixed-domain training achieves 91% on Graph Coloring. For scalability, we use a two-stage approach: a GBDT filter (chen2016xgboost) reduces the action space to 40 candidates, then the GAT policy selects from this set, providing 10–20 speedup while maintaining 95% performance.
@@ Appendix E Additional Experimental Results
Table 9 provides detailed per-domain results for the MiniZinc Challenge dataset, broken down by problem category (van2011minizinc). MetaJuLS achieves consistent improvements across all categories, with particularly strong performance on scheduling and routing problems where the learned policies can exploit structural patterns (perron2009constraint; apt2003principles).
@@ Appendix F Policy Visualization
Figure 4 shows a detailed visualization of the learned policy’s decision-making process on a representative constraint-based instance (velickovic2018graph). The figure illustrates how the GAT attention weights evolve over the course of inference, with the policy initially focusing on constraints that strongly couple many variables and gradually shifting attention to more local constraints as global structure stabilizes (vaswani2017attention; schulman2017proximal).
This visualization demonstrates that the learned policy exhibits sophisticated temporal reasoning, adapting its focus based on the current solver state rather than applying a static heuristic throughout the search process (sutton2018reinforcement; schulman2017proximal).

### v3

@@ Appendix B Mechanistic Analysis and Policy Interpretability
We also analyze attention patterns conditioned on linguistic structures. Constraints involving coordination receive 1.8 higher attention than those involving simple modification, and the policy learns to delay propagation of constraints in garden-path regions until sufficient context is available, mimicking human parsing strategies (Lewis and Steedman, 2014).
On sentences with nested relative clauses (e.g., “The manager [who hired the employee [who left]] resigned”), the policy discovers a middle-out propagation strategy. Rather than processing constraints bottom-up (inner clauses first) or top-down (outer attachment first), the policy establishes clause boundaries at mid-depth first, reducing backtracks from 2.3 to 1.5 on average compared to standard heuristics. This strategy emerges without explicit supervision, suggesting the policy discovers structural regularities through reward optimization alone Shihab et al. (2026). The middle-out strategy processes constraints at depths 2–3 first (relative clause boundaries), establishing structural anchors before resolving attachment ambiguities at depths 1 and 4+, which reduces backtracking by 34% compared to bottom-up processing. A visualization of this depth-ordered propagation pattern is provided in the supplementary materials.
@@ B.1 Linguistic Error Analysis
We characterize error cases by sentence length, ambiguity, constraint density, and schema depth. Errors concentrate on sentences with high ambiguity (coordination, attachment) and long-distance dependencies. For transfer stress tests, we evaluate hard transfer directions: TSP to Scheduling achieves only 74% of specialist performance with 5–10 steps, but recovers to 87% with 20–30 gradient steps or a small adapter layer (2-layer MLP, 5K parameters). To test robustness to determinism violations, we simulate stochastic propagators (randomized pruning order with 10% noise) and find that the policy maintains 92% of deterministic performance, with entropy-aware fallback Shihab et al. (2025) providing additional robustness.
@@ Appendix C Implementation Details
This appendix provides additional implementation details for reproducibility. The MetaJuLS solver is implemented in Python 3.8 using PyTorch 1.12 for the neural network components (LeCun et al., 2015; Goodfellow et al., 2016) and the OR-Tools library for the base CP solver functionality (Perron and Furnon, 2024; Schulte et al., 2010). The GAT architecture uses the PyTorch Geometric library for efficient graph operations (Veličković et al., 2018; Kipf and Welling, 2017).
The training procedure uses a distributed setup with 8 GPU workers, each generating episodes in parallel (Schulman et al., 2017; Lillicrap et al., 2015). For each training round, workers collect a fixed number of episodes, which are then aggregated and used for several epochs of PPO updates before being discarded, following the standard on-policy protocol without long-term replay (Mnih et al., 2015; Silver et al., 2016).
@@ C.1 Hyperparameters and Robustness
| Optimizer | Adam (Kingma and Ba, 2014) |
@@
Table 8 details the full model configuration. The reward function hyperparameters are set to and based on preliminary experiments that balanced domain reduction and the deterministic cost proxy described in the main text (Sutton and Barto, 2018; Schulman et al., 2017). The discount factor encourages long-term planning while preventing excessive focus on distant future rewards (Schulman et al., 2017; Silver et al., 2017).
@@ Appendix D Constraint Programming Experiments: Extended Results
Having established MetaJuLS’s effectiveness on linguistic constraints, we test universality: do the learned scheduling principles transfer to non-linguistic constraint satisfaction? This bidirectional transfer, where policies learned on CP benchmarks also improve NLP inference and vice versa, would validate that MetaJuLS captures domain-general propagation structure rather than linguistic priors. We evaluate on classical constraint programming benchmarks from MiniZinc and XCSP (Nethercote et al., 2011; Lecoutre et al., 2011), training on 500 Knapsack, 300 TSP, and 200 Graph Coloring instances. We compare against JuLS (FIFO), random scheduling, dom/wdeg heuristics, Google OR-Tools (Perron and Furnon, 2024), and Activity-Based adaptive heuristics (Marques-Silva et al., 2012), using identical time limits (1200s per instance) and hardware.
@@
On the XCSP Competition dataset, which focuses on high-difficulty instances (Lecoutre et al., 2011), the performance advantages of MetaJuLS are even more pronounced. The learned policies excel at identifying bottleneck constraints in complex constraint graphs (Veličković et al., 2018), leading to more effective search space pruning (Bessière, 2006). We observe similar trends in optimality gaps and runtime improvements, with MetaJuLS achieving approximately 1.5 speedup over OR-Tools on average (Perron and Furnon, 2024).
@@ D.1 Policy Analysis and Transfer
The learned policies reveal both rediscovery of classical heuristics and novel strategies. On Knapsack instances, the policy correlates strongly () with the Domain-over-Weight heuristic (Apt, 2003), validating that learning recovers known effective strategies. On Graph Coloring, the policy learns to focus on vertices with high violation potentials rather than degree alone, outperforming classical heuristics on dense graphs (Régin, 1994; Gent et al., 2006). Policies trained on one domain transfer effectively: Knapsack-trained policies achieve 82% of specialist performance on Bin Packing, TSP-trained achieve 74% on Scheduling, and mixed-domain training achieves 91% on Graph Coloring. For scalability, we use a two-stage approach: a GBDT filter (Chen and Guestrin, 2016) reduces the action space to 40 candidates, then the GAT policy selects from this set, providing 10–20 speedup while maintaining 95% performance.
@@ Appendix E Additional Experimental Results
Table 9 provides detailed per-domain results for the MiniZinc Challenge dataset, broken down by problem category (Nethercote et al., 2011). MetaJuLS achieves consistent improvements across all categories, with particularly strong performance on scheduling and routing problems where the learned policies can exploit structural patterns (Perron et al., 2009; Apt, 2003).
@@ Appendix F Policy Visualization
Figure 4 shows a detailed visualization of the learned policy’s decision-making process on a representative constraint-based instance (Veličković et al., 2018). The figure illustrates how the GAT attention weights evolve over the course of inference, with the policy initially focusing on constraints that strongly couple many variables and gradually shifting attention to more local constraints as global structure stabilizes (Vaswani et al., 2017; Schulman et al., 2017).